OpenStudy (cali2000):

I need help, I am so confused. I wish I understood these things better. solve the system by substitution -4.5x - 2y = -12.5 3.25x - y = -0.75

OpenStudy (studygurl14):

Notice how the second equation has -y. The easiest thing would be to rearrange the second equation so that y is by itself on the left side

OpenStudy (cali2000):

so i would subtract 3.25x to each side right?

OpenStudy (studygurl14):

yes

OpenStudy (radar):

A good plan,

OpenStudy (cali2000):

so it would be -y = -0.75 - 3.25x ??

OpenStudy (studygurl14):

Perfect. now get rid of the negative

OpenStudy (studygurl14):

on the left side

OpenStudy (cali2000):

so where would I go from here, I am so bad at solving these I'm sorry

OpenStudy (studygurl14):

divide both sides by -1

OpenStudy (cali2000):

where did -1 come from?

OpenStudy (studygurl14):

-y is basically just (-1)y

OpenStudy (cali2000):

OH yeah

OpenStudy (cali2000):

y = .75 - 3.25x

OpenStudy (studygurl14):

almost. y = 0.75 + 3.25x Now plug that into the first equation for y

OpenStudy (studygurl14):

and solve for x

OpenStudy (studygurl14):

That's the substitution part

OpenStudy (cali2000):

-4.5 (3.25) - 2 (.75) = -12.5

OpenStudy (cali2000):

i dont think i arranged that correctly

OpenStudy (studygurl14):

Not quite. -4.5x -2(0.75 + 3.25) = -12.5

OpenStudy (studygurl14):

sorry 3.25x I meant

OpenStudy (studygurl14):

-4.5 -2(0.75 + 3.25x) = -12.5 solve for x

OpenStudy (cali2000):

i've been trying to figure that out and i can't get it, do i multiply??

OpenStudy (radar):

Remove the parentheses getting -4.5 - 1.5 - 6.5x = -12.5 Now i bet you can solve for x.

OpenStudy (radar):

Add a +6 to both sides getting -6.5x = 6.5 Now divide both sides by a -6.5 so that we have a positive x on the left or: x = -1 Now your work is not over you still have to solve for y. Substitute the -1 for x in the second original equation: 3.25x - y = -0.75 and solve for y.